A calendar-year taxpayer acquired four new machines in 2019 on the dates shown below. Section 179 is not elected, but regular MACRS is used to depreciate the machines. The machines have a recovery period of 7 years. The taxpayer elects out of taking bonus depreciation.

February 1 $50,000

April 1 70,000

October 1 60,000

December 1 80,000

a. compute total MACRS depreciation expense for 2019 that would be reported on Schedule C, using the appropriate averaging convention.

b. Same as in part A., except that the third machine was acquired on September 30 (instead of October 1) Comment on the significance of the difference between this amount and your answer to part a.
Answer #1

a) Depreciation = (Purchased cost - salvage value)/ Life of asset

Life of the machine1 is 7 years and salvage value is nill

Depreciation = 50000/ 7 years = 7142.86

As the machinery 1 was purchased on feb1 depreciation will be charged for 11 months

=(7142.86/12)*11 = 6,548

* Machinery 2 is purchased on 1 April, therefore its depreciation will be calculated for 9 months

=((70000/7 years)/12)*9 months

= 7,500

* Machinery 3 is purchased on 1 October, therefore its depreciation will be calculated for 3 months

=((60000/7years)/12)*3 months

=2,143

* Machinery 4 is purchased on 1 December, therefore its depreciation will be calculated for 1 months

=((80000/7 years)/12)*1 month

=952

Therefore, total depreciation to be reported in schedule C = 6548 + 7500 + 2143 + 952

= 17,143

b) Here third machine is considered as aquired on 30 september instead of 1 october, which means that there will be 1 more additional day of use of machine in the year

Total days in each month for current year is calculated as follows:

September = 1 day

October = 31 days

November = 30 days

December = 31 days

Total = 93 days

Therefore depreciation will be calcualted for machinery 3 for 93 days as follows:

= (8,571.43/3650*93 days = 2,188

Total depreciation reported in schedule C will be = 6548 + 7500 + 2188 + 952 = 17,188

Therefore total difference between pat a and part b = 17,188 - 17,143 = $45.